# Algorithms for Generating Small Random Samples

### v2

@@ Abstract
This report presents algorithms for generating small random samples without replacement. It considers two cases. It presents an algorithm for sampling a pair of distinct integers, and an algorithm for sampling a triple of distinct integers. The worstcase runtime of both algorithms is constant, while the worstcase runtime of common algorithms for the general case of sampling $k$ elements from a set of $n$ are linear in $n$. Java implementations of both algorithms are included in the open source library $\rho\mu$.
@@
algorithm; random sampling; Java; open source
@@ 1 Introduction
Efficiently generating random samples of $k$ elements from a set of $n$ elements, without replacement, is important to a variety of applications. There are algorithms for handling this general case, such as reservoir sampling [11], pool sampling [5], and insertion sampling [2]. Knuth [6] discusses a variety of algorithms as well. The runtime of both reservoir sampling and pool sampling is $O(n)$. While generating the sample, reservoir sampling requires $(n-k)$ random integers, many more than the sample size $k$ when $k$ is small. It is a great choice when $k$ is large, however. Pool sampling requires $k$ random numbers to generate a sample of size $k$, which is good when $k$ is small, however its runtime grows linearly in $n$; and pool sampling also requires $O(n)$ extra memory. Insertion sampling was designed with small samples in mind, with a runtime of $O(k^{2})$, requiring $k$ random numbers to generate a sample of size $k$. It is an efficient choice when you know $k$ will be small, although for large $k$ it is much slower than the reservoir and pool sampling algorithms.
@@
Although the insertion sampling algorithm handles small $k$ nicely, we can do much better with algorithms designed specially for the specific value of $k$. In this report, we present algorithms for the special cases of $k=2$ and $k=3$. The runtime of both algorithms is $O(1)$, with the algorithm for generating a random pair of distinct integers requiring 2 random numbers, and the algorithm for generating a random triple of distinct integers requiring 3 random numbers. Our original motivation for efficient algorithms for randomly sampling pairs and triples of distinct integers is in the implementation of various mutation and crossover operators for evolutionary algorithms for the space of permutations [4], where such evolutionary operators require generating random combinations of indexes. In such an application, one needs to generate random samples from the integer interval $[0,n)$. Thus, the algorithms that we present are specified as sampling from that interval. However, without loss of generality, both algorithms are applicable for the broader case of sampling pairs and triples of elements from a set of $n$ elements, provided you define a mapping of the $n$ elements to the integers in $[0,n)$.
@@
The two algorithms for sampling pairs and triples of integers are presented in Section 2. We provide Java implementations in the open source library $\rho\mu$, which provides a variety of randomization-related utilities [3]. We empirically compare the runtime performance of the algorithms compared to existing random sampling algorithms using the $\rho\mu$ library, describing our experimental methodology in Section 3 and the results in Section 4. We wrap up in Section 5.
@@ 2 Algorithms
Algorithm 1, $\mathrm{RandomPair}(n)$, generates a random pair $(i,j)$ of distinct integers from the interval $[0,n)$, uniformly distributed over the space of all ($n\choose 2$ such pairs. Line 1 generates the first of the two integers, $i$, uniformly from the interval $[0,n)$. Then, to generate the other integer, a random integer $j$ uniformly distributed over the interval $[0,n-1)$ is first generated (line 2). If $j$ is the same as $i$, we reset $j$ to $n-1$ (lines 3–4), since $n-1$ was excluded via the half-open interval $[0,n-1)$.
@@
3:  if $j=i$ then
4:        $j\leftarrow n-1$
5:  return  $(i,j)$
Algorithm 2, $\mathrm{RandomTriple}(n)$, generates a random triple $(i,j,k)$ of distinct integers from the interval $[0,n)$, uniformly distributed over the space of all ($n\choose 3$ such pairs. It initially generates $i$, $j$, and $k$ uniformly at random from the half-open intervals $[0,n)$, $[0,n-1)$, and $[0,n-2)$, respectively. The remainder of the algorithm handles mapping duplicates to the values excluded by the half-open intervals. The adjustment in lines 4–5 handle the case when $i$ and $j$ are identical in the same way as in Algorithm 1, mapping the duplicate to $n-1$. Lines 6–7 maps $k$ to $n-2$ if it is initially equal to $i$. It can safely do this without comparing to $j$ since $j$ had just been set to $n-1$. Lines 9–12 handle the case when $i$ and $j$ are already initially different. In that case, we must compare $k$ to both $i$ and $j$ in sequence, mapping $k$ to one of the two values excluded by its half-open interval if it is the same as either $i$ or $j$. Both comparisons are always necessary since the adjustment on line 10 could potentially create a conflict between $k$ and $i$ that must then be resolved.
@@
4:  if $j=i$ then
5:        $j\leftarrow n-1$
6:        if $k=i$ then
7:              $k\leftarrow n-2$
8:  else
9:        if $k=j$ then
10:              $k\leftarrow n-2$
11:        if $k=i$ then
12:              $k\leftarrow n-1$
13:  return  $(i,j,k)$
The worstcase runtime of both algorithms is $O(1)$. Algorithm 1 and Algorithm 2 generate 2 and 3 random integers, respectively; and both have constant numbers of comparisons and assignments to adjust for duplicates.
@@ 3 Experimental Methodology
We implemented the new algorithms from Section 2, as well as other random sampling algorithms in the open source library $\rho\mu$. The experiments of this paper specifically use $\rho\mu$ 4.0.0. The code to reproduce the experiments is also available on GitHub, as is the raw data from my runs of the experiments. See Table 1 for URLs to the libary sourcecode, project website, and in the Maven Central repository, as well as for the source code and raw data from the experiments.
@@
Table 2 lists the methods of $\rho\mu$’s EnhancedRandomGenerator class that are used in the experiments. The EnhancedRandomGenerator class wraps any class that implements Java 17’s RandomGenerator interface adding a variety of functionality to it, or in some cases replacing implementations with more efficient algorithms. There are multiple nextIntPair and nextIntTriple methods, which differ in terms of return type (e.g., returning an array of type int versus returning library specific record IndexPair and IndexTriple).
@@
insertion sampling [2] sampleInsertion
pool sampling [5] samplePool
reservoir sampling [11] sampleReservoir
We use OpenJDK 64-Bit Server VM version 17.0.2 in the experiments on a Windows 10 PC with an AMD A10-5700, 3.4 GHz processor and 8GB memory. For the pseudorandom number generator (PRNG), we use Java’s SplittableRandom class, which implements the SplitMix [10] algorithm, which is a faster optimized version of the DotMix [7] algorithm, and which passes the DieHarder [1] tests.
The $\rho\mu$ library’s EnhancedRandomGenerator class replaces Java’s method for generating random integers subject to a bound with an implementation of a much faster algorithm [8]. Our prior experiments [3] show that Lemire’s approach uses less than half the CPU time as Java’s RandomGenerator.nextInt(bound) method. So although we are using Java’s SplittableRandom class as the PRNG, by wrapping it in an instance of $\rho\mu$’s EnhancedRandomGenerator, we are significantly speeding the runtime of all of the algorithms in our experiments.
We use the Java Microbenchmark Harness (JMH) [9] to implement our experiments. For each experiment condition (e.g., algorithm and value of $n$) we use five 10-second warmup iterations to ensure that the Java JVM is properly warmed up, and we likewise use five 10-second iterations for measurement. We measure and report average time per operation in nanoseconds, along with 99.9% confidence intervals.

### v3

@@ Abstract
[Abstract]We present algorithms for generating small random samples without replacement. We consider two cases. We present an algorithm for sampling a pair of distinct integers, and an algorithm for sampling a triple of distinct integers. The worst-case runtime of both algorithms is constant, while the worst-case runtimes of common algorithms for the general case of sampling $k$ elements from a set of $n$ increase with $n$. Java implementations of both algorithms are included in the open source library $\rho\mu$.
@@
algorithm, random sampling, Java, open source, small samples
articletype: Short Communicationjournal: Software: Practice and Experiencevolume: 00footnotetext: Abbreviations: JMH, Java microbenchmark harness; JVM, Java virtual machine; PRNG, pseudorandom number generator.
@@ 1 Introduction
Efficiently generating random samples of $k$ elements from a set of $n$ elements, without replacement, is important to a variety of applications. There are algorithms for handling this general case, such as the reservoir sampling family of algorithms 1, 2, pool sampling 3, insertion sampling 4, among others 5, 6. The runtime of pool sampling and some forms of reservoir sampling is $O(n)$. Other forms of reservoir sampling 2 improve the runtime to $O(k(1+\ln(n/k)))$, but this still increases with $n$. Insertion sampling was designed with small samples in mind, with a runtime of $O(k^{2})$, requiring $k$ random numbers to generate a sample of size $k$. It is an efficient choice when you know $k$ will be small, although for large $k$ it is much slower than alternatives. Others such as Ting make other improvements to the classical random sampling algorithms 7. There are also several parallel algorithms for random sampling without replacement 8, weighted random sampling 9, and for random sampling (both weighted and unweighted) from data streams 10, 11, 12.
Although insertion sampling handles small $k$ nicely, we can do much better with algorithms designed for the specific value of $k$. For example, the special case of $k=1$ has been widely studied, such as Lemire’s algorithm for random integers in an interval that is significantly faster than the alternatives provided within common programming languages by avoiding nearly all division operations 13, or Goualard’s approach to sampling random floating-point numbers from an interval 14. It is common for applications to require repeatedly generating random integers from an interval, such as algorithms for shuffling arrays, as well as sampling algorithms for the general case. Brackett-Rozinsky and Lemire propose an algorithm for sampling multiple independent bounded random integers from a single random word, and demonstrate how the technique can speed up algorithms for shuffling 15.
@@
In this article, we present algorithms for the special cases of $k=2$ and $k=3$. The runtime of both algorithms is $O(1)$, with the algorithm for generating a random pair of distinct integers requiring two random numbers, and the algorithm for generating a random triple of distinct integers requiring three random numbers. Our original motivation for efficient algorithms for randomly sampling pairs and triples of distinct integers is in the implementation of various mutation and crossover operators for evolutionary algorithms for the space of permutations 16, where such evolutionary operators require generating random combinations of indexes. In such an application, one requires random samples from the integer interval $[0,n)$. Thus, the algorithms that we present are specified as sampling from that interval. However, without loss of generality, both algorithms are applicable for the broader case of sampling pairs and triples of elements from a set of $n$ elements, provided you define a mapping of the $n$ elements to the integers in $[0,n)$. We utilize Lemire’s algorithm for bounded random integers 13 within our implementation of the algorithms of this paper, as well as in our implementations of the general sampling algorithms used in the experiments. The Brackett-Rozinsky and Lemire approach to multiple bounded random integers 15 can potentially enable further optimizations as well, but we have not explored doing so.
@@
The two algorithms for sampling pairs and triples of integers are presented in Section 2. We also discuss in that section how one can derive a special purpose sampling algorithm for any fixed $k$ in the form of a sampling network, although for larger $k$ this is likely impractical, and demonstrate with an algorithm for the case of $k=4$. We provide Java implementations in the open source library $\rho\mu$, which provides a variety of randomization-related utilities 17. We empirically compare the runtime performance of the algorithms to existing random sampling algorithms using the $\rho\mu$ library, describing our experimental methodology in Section 3 and the results in Section 4. We wrap up in Section 5.
@@ 2 Algorithms
Algorithm 1, $\mathrm{RandomPair}(n)$, generates a random pair $(i,j)$ of distinct integers from the interval $[0,n)$, uniformly distributed over the space of all $2{n\choose 2}$ distinct pairs. Line 1 generates the first of the two integers, $i$, uniformly from the interval $[0,n)$. Then, a random integer $j$ uniformly distributed over the interval $[0,n-1)$ is generated (line 2); and if $j$ is the same as $i$, we reset $j$ to $n-1$ (line 3), since $n-1$ was excluded via the half-open interval $[0,n-1)$.
@@
3:if $j=i$ then $j\leftarrow n-1$
4:return $(i,j)$
Algorithm 2, $\mathrm{RandomTriple}(n)$, generates a random triple $(i,j,k)$ of distinct integers from the interval $[0,n)$, uniformly distributed over the space of all $6{n\choose 3}$ distinct triples. It initially generates $i$, $j$, and $k$ uniformly at random from the half-open intervals $[0,n)$, $[0,n-1)$, and $[0,n-2)$, respectively. The remainder of the algorithm maps duplicates to the values excluded by the half-open intervals. Line 4 ensures that $j$ and $k$ are different. Lines 5–6 in turn adjusts if necessary so that neither are the same as $i$.
@@
4:if $k=j$ then $k\leftarrow n-2$
5:if $j=i$ then $j\leftarrow n-1$
6:if $k=i$ then $k\leftarrow n-1$
7:return $(i,j,k)$
The worst-case runtime of both algorithms is $O(1)$. Algorithm 1 and Algorithm 2 generate 2 and 3 random integers, respectively; and both have constant numbers of comparisons and assignments to adjust for duplicates.
The bounds of the random numbers utilized by Algorithm 1 and Algorithm 2 are similar to those of the first two and three steps of a Fisher-Yates shuffle 18 as well as the first two to three iterations of insertion sampling (see later in Section 3.1). Additionally, each if-statement that resolves duplicates is reminiscent of the “compare-exchange” operation of sorting networks 19. But, instead of a “compare-exchange,” it is a “compare-change.” This suggests the potential to derive a sampling network from the approach. Such a sampling network may have similar benefits as sorting networks, such as lending themselves well to hardware implementation and parallelism. For example, the execution order of lines 5 and 6 of Algorithm 2 does not matter. It is straightforward to derive a corresponding sampling network for any fixed $k$, although the benefit in software diminishes since the obvious approach leads to a number of compare-changes that increases quadratically in $k$. For example, Algorithm 3 shows how to extend the approach to $k=4$. A hypothetical circuit implementation could align the compare-changes in layers of independent operations, such as lines 6–7 in parallel and lines 8–10 in parallel.
Algorithm 3 $\mathrm{RandomFourTuple}(n)$
1:$h\leftarrow\mathrm{Rand}(n)$
2:$i\leftarrow\mathrm{Rand}(n-1)$
3:$j\leftarrow\mathrm{Rand}(n-2)$
4:$k\leftarrow\mathrm{Rand}(n-3)$
5:if $k=j$ then $k\leftarrow n-3$
6:if $j=i$ then $j\leftarrow n-2$
7:if $k=i$ then $k\leftarrow n-2$
8:if $i=h$ then $i\leftarrow n-1$
9:if $j=h$ then $j\leftarrow n-1$
10:if $k=h$ then $k\leftarrow n-1$
11:return $(h,i,j,k)$
@@ 3 Experimental Methodology
3.1 General Sampling Algorithms
We experimentally validate the approach relative to existing algorithms for the general case of sampling $k$ distinct elements from a set of $n$ to explore the benefit of special purpose algorithms when $k$ is known and small.
First, we consider two forms of reservoir sampling 1, 2, which refers to a family of sampling algorithms. Reservoir algorithms were originally introduced decades ago for the problem of randomly sampling $k$ records from an unknown number of records $n$ via a single pass over the records. But they are also easily adapted to the simpler problem of sampling integers with known $n$. The first reservoir sampling algorithm that we consider is the simplest, and referred to by Vitter as algorithm R 1. We provide pseudocode in Algorithm 4. It begins by initializing the sample with the first $k$ elements (lines 1–2). It then iterates over the remaining $n-k$ elements in a very similar way to an array shuffling procedure, choosing a random index $j$ for the next element (lines 3–4). If that index is in the bounds of the sample, that element replaces the one at that index in the sample (line 5). The runtime of this reservoir sampling algorithm is $O(n)$, and it requires $O(n-k)$ random numbers, which is desirable when $k$ is large, but there are better algorithms for small $k$. In the same article, Vitter described a series of improvements leading to his algorithms X, Y, and Z, the last of which having an expected runtime of $O(k(1+\ln(n/k)))$.
Algorithm 4 $\mathrm{ReservoirSamplingR}(n,k)$
1:samplea new array of $\mathrm{sample}\leftarrow\text{a new array of length}\ k$
2:for $i=0$ to $k-1$ do $\mathrm{sample}[i]\leftarrow i$
3:for $i=k$ to $n-1$ do
4:     $j\leftarrow\mathrm{Rand}(i+1)$
5:     if $j<k$ then $\mathrm{sample}[j]\leftarrow i$
6:end for
7:return $\mathrm{sample}$
The other reservoir sampling algorithm that we consider is Li’s algorithm L 2, which is simpler and an improvement over Vitter’s algorithm Z 1. We provide pseudocode for the L form of reservoir sampling in Algorithm 5. In the pseudocode, $U()$ returns a random floating-point value in $[0.0,1.0)$. It begins like the previous reservoir algorithm by initializing the sample with the first $k$ elements (lines 1–2). Instead of explicitly iterating over all $n-k$ remaining elements, it skips groups of elements. See Li for derivation of the skip lengths and the details of the runtime analysis 2. The expected runtime still grows with $n$, but much better than linearly, $O(k(1+\ln(n/k)))$. As $k$ approaches $n$, this converges to $O(k)$. Thus, this is an especially good choice for larger $k$. There are $k$ steps to initialize the sample, and the expected number of iterations of the while loop is $O(k\ln(n/k))$. It should be noted that the iterations are more costly with each iteration generating two uniform random floating-point values and a random bounded integer, along with an exponentiation and three logs.
Algorithm 5 $\mathrm{ReservoirSamplingL}(n,k)$
1:samplea new array of $\mathrm{sample}\leftarrow\text{a new array of length}\ k$
2:for $i=0$ to $k-1$ do $\mathrm{sample}[i]\leftarrow i$
3:$w\leftarrow\exp\left(\frac{\ln(U())}{k}\right)$
4:$i\leftarrow k+\left\lfloor\frac{\ln(U())}{\ln(1-w)}\right\rfloor$
5:while i < n do
6:     $\mathrm{sample}[\mathrm{Rand}(k)]\leftarrow i$
7:     $w\leftarrow w\cdot\exp\left(\frac{\ln(U())}{k}\right)$
8:     $i\leftarrow i+1+\left\lfloor\frac{\ln(U())}{\ln(1-w)}\right\rfloor$
9:end while
10:return $\mathrm{sample}$
Next consider pool sampling 3, which has additional names in the literature, including random draw sampling 20. Its authors originally referred to it generically as “SELECT” 3, but there are others with that same name. Algorithm 6 shows it in pseudocode form. It maintains a pool of unused elements (initialized in lines 2–4 and updated in lines 8–9), and draws elements randomly from the pool (lines 6–7) until the sample is complete. Pool sampling requires $k$ random numbers to generate a sample of size $k$, which is good when $k$ is small, however its runtime grows linearly in $n$; and it also requires $O(n)$ extra storage. Ernvall and Nevalainen 5 use a hash table to reduce the extra storage to $O(k)$ and the runtime to $O(k^{2})$.
Algorithm 6 $\mathrm{PoolSampling}(n,k)$
1:samplea new array of $\mathrm{sample}\leftarrow\text{a new array of length}\ k$
2:poola new array of $\mathrm{pool}\leftarrow\text{a new array of length}\ n$
3:for $i=0$ to $n-1$ do $\mathrm{pool}[i]\leftarrow i$
4:$\mathrm{remaining}\leftarrow n$
5:for $i=0$ to $k-1$ do
6:     $j\leftarrow\mathrm{Rand}(\mathrm{remaining})$
7:     $\mathrm{sample}[i]\leftarrow\mathrm{pool}[j]$
8:     $\mathrm{remaining}\leftarrow\mathrm{remaining}-1$
9:     $\mathrm{pool}[j]\leftarrow\mathrm{pool}[\mathrm{remaining}]$
10:end for
11:return $\mathrm{sample}$
Like pool sampling, the insertion sampling 4 algorithm requires only $k$ random numbers. But unlike pool sampling, its storage requirement is constant. Its runtime is $O(k^{2})$. It is not widely studied as it was introduced within the context of a specific application, one where $k$ is unknown ahead of time but likely small. Algorithm 7 provides the details. It maintains the currently sampled elements in sorted order (lines 5, 7–10). It iteratively generates a sequence of random integers $v\in[\mathrm{Rand}(n),\mathrm{Rand}(n-1),\ldots,\mathrm{Rand}(n-k+1)]$ (line 3). Each $v$ is treated as an index into an ordered (implicit) list of the unused elements. Line 6 skips over used elements as $v$ is inserted into the sample in its sorted position (line 10).
Algorithm 7 $\mathrm{InsertionSampling}(n,k)$
1:samplea new array of $\mathrm{sample}\leftarrow\text{a new array of length}\ k$
2:for $i=0$ to $k-1$ do
3:     $v\leftarrow\mathrm{Rand}(n-i)$
4:     $j\leftarrow k-i$
5:     while $j<k$ and $v\geq\mathrm{sample}[j]$ do
6:         $v\leftarrow v+1$
7:         $\mathrm{sample}[j-1]\leftarrow\mathrm{sample}[j]$
8:         $j\leftarrow j+1$
9:     end while
10:     $\mathrm{sample}[j-1]\leftarrow v$
11:end for
12:return $\mathrm{sample}$
The proposed algorithms in Section 2 (Algorithms 12, and 3) produce samples that are not only uniformly random over combinations, but also uniformly random over permutations. This is also true of pool sampling (Algorithm 6). However, it is not true of the other reference algorithms. Insertion sampling (Algorithm 7) generates an ordered sample. For uniform ordering, shuffle the sample before the return on line 12. Line 2 of both versions of reservoir sampling (Algorithms 4 and 5) constrains the possible positions of the first $k$ elements. Shuffling the initial sample of line 2 achieves uniform random ordering. In all three cases, the adjustment introduces an $O(k)$ step with $O(k)$ additional random numbers. In the experiments, we use these reference algorithms as specified in the pseudocode and do not introduce such shuffling steps.
3.2 Methodology
@@
We implemented the new algorithms from Section 2, and three of the other sampling algorithms from Section 3.1 in the open source library $\rho\mu$. The experiments use $\rho\mu$ 4.2.0. The code for reservoir (L) sampling is available in the GitHub repository that contains the code to reproduce the experiments, and the raw data from my runs. See Table 1 for URLs to the library source code, project website, and in the Maven Central repository, as well as for the source code and raw data from the experiments.
@@
Table 2 lists the methods of $\rho\mu$’s EnhancedRandomGenerator class that are used in the experiments. The EnhancedRandomGenerator class wraps any class that implements Java 17’s RandomGenerator interface adding a variety of functionality to it, or in some cases replacing implementations with more efficient algorithms. There are multiple nextIntPair and nextIntTriple methods, which differ in terms of return type (e.g., returning an array of type int versus returning a library specific record IndexPair and IndexTriple).
@@
insertion sampling 4 sampleInsertion
pool sampling 3 samplePool
reservoir (R) sampling 1 sampleReservoir
reservoir (L) sampling 2 ReservoirL class in experiment repository
We use OpenJDK 64-Bit Server VM version 17.0.2 in the experiments on a Windows 10 PC with an AMD A10-5700, 3.4 GHz processor and 8GB memory. For the pseudorandom number generator (PRNG), we use Java’s SplittableRandom class, which implements the SplitMix 21 algorithm, which is a faster optimized version of the DotMix 22 algorithm, and which passes the DieHarder 23 tests.
The $\rho\mu$ library’s EnhancedRandomGenerator class replaces Java’s method for generating random integers subject to a bound with an implementation of a much faster algorithm 13. Our prior experiments 17 show that Lemire’s approach uses less than half the CPU time as Java’s RandomGenerator.nextInt(bound) method. So although we are using Java’s SplittableRandom class as the PRNG, by wrapping it in an instance of $\rho\mu$’s EnhancedRandomGenerator, we are significantly speeding the runtime of all of the algorithms in our experiments.
We use the Java Microbenchmark Harness (JMH) 24 to implement our experiments. JMH is developed by the same team that develops the OpenJDK, and is the preferred harness for benchmarking Java code. Since the Java Virtual Machine (JVM) is adaptive, utilizing just-in-time compilation of hot-spots identified at runtime to boost performance, it is important to warm up the JVM to achieve a steady-state before benchmarking. The JMH handles this warmup phase, and other Java benchmarking related challenges. For each experiment condition (e.g., algorithm and value of $n$) we use five 10-second warmup iterations to ensure that the Java JVM is properly warmed up, and we likewise use five 10-second iterations for measurement. JMH executes the benchmark method as many times as each 10-second iteration allows, which leads to hundreds of millions of method invocations for each experiment condition. We use JMH’s built-in blackhole to consume the generated samples to prevent dead-code elimination. We measure and report average time per operation in nanoseconds, along with 99.9% confidence intervals as calculated by JMH. The times reported by JMH are not CPU times. Although not indicated in the JMH documentation, examination of the JMH source code 24 reveals that it calculates elapsed time using the JVM’s high-resolution time source via the System.nanoTime() method, which returns nanoseconds since an arbitrary origin. We round the averages to the nearest tenth of a nanosecond since the JVM’s time source is of nanosecond precision (JMH’s reported averages are to thousandths of a nanosecond, a false level of precision). A consequence of elapsed time is that benchmarks can be affected by the state of the system (e.g., competition for system resources). To minimize such impact, we shut down all applications under our control, and disabled various non-critical background processes (e.g., auto-updaters, etc) to limit background processes as much as feasible.